The Cobb- Douglas production function is given by the form Y = A La Kb

Where, Y = Output (mt/day), L = Labour (hours/mt) K = Capita (Rs/mt). By taking the natural logarithm of each term in the function, that is LnY = LnA+aLnL+bLnK. Using a standard multiple linear regression model the parameters of the function were obtained as follows;

Ln Y = 4.45 +0.5LnL + 0.4LnK

(2.2) (0.01) (0.02)

The standard error for each estimated parameter is shown in parentheses below the estimate.

Construct equation for Cobb- Douglas production function and briefly explain the properties of this production function (4 marks)

Measure marginal product of labour and average product of capital when L=10 and K=5 (3 marks)

Is this production function exhibit the law of diminishing returns in the short run? (3 marks)


1
Expert's answer
2021-09-07T08:24:03-0400

Ln Y = 4.45 +0.5LnL + 0.4LnK

Differentiating both sides w.r.t. labour we get:

"\frac{1}{Y}""\times""\frac{dY}{DL}=""\frac{0.5}{L}"


"\frac{dY}{dL}=""\frac{0.5Y}{L}"

Elasticity of labour "=""\frac{dY}{DL}""\times""\frac{L}{Y}"


E(L)="\frac{0.5Y}{L}" "\times" "\frac{L}{Y}"

E(L)=0.5


Elasticity of capital ="\frac{dY}{dK}\times\frac{K}{Y}"

Differentiating output equation w.r.t to capital we get

"\frac{1}{Y}\times\frac{dY}{dK}=\frac{0.4}{K}"

Putting in elasticity equation we get;

E(K)="\frac{0.4Y}{K}\times\frac{K}{Y}"

E(K)=0.4

Therefore, cobb douglas function is given by

Q = 4.45L0.5K0.4


properties of Cobb-Douglas production function

  1. Homogeneous of Degree One i.e Q=ALαKβ. α + β= 1
  2. Average and marginal products of L and K, i.e., APL, APK, MPL, and MPK would all be the functions of L-K or K-L ratio.
  3. MRTSL.K = "\frac{MPL}{MPK}"  = function of "\frac{L}{K}" ratio

marginal product of labour

MPL is calculated by partially differentiating the cobb douglas function w.r.t. to L

MPL= 2.225 L-0.5K0.4

But L=10 and K= 5

Therefore MPL = 1.354

average product of capital

APK= "\frac{Q}{K}" = ALαKβ-1

APK= 4.45L0.5K-0.6

But L=10 and K= 5

APK = 4.45 * 100.5*5-0.6

APK = 5.36


This function exhibits the law of diminishing returns in the short run since the variables of labor and capital are varied in isolation.